SAT scores have a mean of 1000 and a standard deviation of 220.

Q18: What is the probability that a random student will score more than 1400?

Q19: Using the Central Limit Theorem, what is the probability that sample of 100 students will have an average score less than 990?

Q20: Using the Central Limit Theorem, what is the probability that sample of 100 students will have an average score between 990 and 1010?

Use excel functions to calculate your answers.
Answer #1

We have the normal distribution parameters here:

mean = 1000

Stdev = 220

Q18. P(X>1400) = P(Z> (1400-1000)/220) = P(Z>1.818) = 0.0345. Use excel formula = 1 - NORMS.DIST(1.818) to get this

Q19. P(Xbar <990) = P(Z< (990-1000)/(220/sqrt(100)) = .3247. Use Excel formula = 1 - NORMS.DIST(-.4546) = .3247 to get this

Q20. P(990<X<1010) = P(-.4546<Z<.4546) = .6752-.3247 = .3505 ( We used Excel formula NORMS.DIST(.4546) - NORMS.DIST(-.4546) to get the answer.